OpenStudy (red_devil822):

Find the diameter of the circle (x - 4)2 + (y + 6)2 = 16. A) 2 B) 4 Eliminate C) 8 D) 16

Nnesha (nnesha):

\[\huge\rm (x-h)^2+(y-k)^2=r^2\] circle equation where (h,k) is the center and r is radius

Nnesha (nnesha):

we know that \[\huge\rm d= r \cdot 2\] d=diameter

OpenStudy (red_devil822):

ok

Nnesha (nnesha):

so what's the radius ?

OpenStudy (red_devil822):

i dont know what to do to be honest

Nnesha (nnesha):

\[ (x - 4)^2 + (y + 6)^2 = 16\] compare this equation with the general equation of a circle

Nnesha (nnesha):

what number is replaced by r^2 ?

OpenStudy (red_devil822):

i figure it out. is it b?

Nnesha (nnesha):

how did you get b?

Nnesha (nnesha):

read my 2nd comment diameter = radius times 2

OpenStudy (red_devil822):

no c